OpenStudy (anonymous):

(a) Estimate the area under the graph of f(x)=7/x from x=1 to x=5 using four approximating rectangles and right endpoints. (b) Repeat part (a) using left endpoints.

OpenStudy (anonymous):

For the right endpoints would it just be \[\frac{ 7 }{ 2 } +\frac{ 7 }{ 3 }+\frac{ 7 }{ 4 }+\frac{ 7 }{ 5 }\] ?

OpenStudy (tommynaut):

That looks right. Using the same mentality, what would part b) be?

OpenStudy (anonymous):

Hmm... I think that's \[7+\frac{ 7 }{ 2 }+\frac{ 7 }{ 3 }+\frac{ 7 }{ 4 }\] am I right?

OpenStudy (zale101):

Correct.

OpenStudy (anonymous):

Thanks so much! :D

OpenStudy (freckles):

the only thing I would add..

OpenStudy (freckles):

or better word multiply by is...

OpenStudy (freckles):

oh the base which is (5-1)/4 which is 1 :) so not really needed

OpenStudy (anonymous):

Ohh right!

OpenStudy (anonymous):

I was too lazy to type that in :P

OpenStudy (anonymous):

but thanks!!!

OpenStudy (freckles):

but you didn't type it in because you knew you didn't need it :)

OpenStudy (anonymous):

True xD

OpenStudy (freckles):

Go you! :)

OpenStudy (anonymous):

Haha thanks @freckles :D